3) Nitrogen gas can be produced by combining gaseous ammonia (NH 3 ) over solid copper (IIoxide at high temperatures. The other products of the reaction are solid copper and water vapor. If a sample containing 35.1 g of NH 3 is reacted with 85.6 g of Cuo. How many grams of will be N 2 formed?


1
Expert's answer
2021-02-15T03:19:39-0500

3CuO + 2 NH3 → 3Cu + N2 + 3H2O

"n(CuO)=\frac{m}{M}=\frac{85.6}{79.545}=1.076"


"n(NH3)=\frac{m}{M}=\frac{35.1}{17.031}=2.061"


calculations are carried out further on the limiting reagent

"\frac{1.076}{3}=\frac{x}{1}"

x=0.359 n(N2)

"m=n\times M=0.359\times28.014=10.05" g